OpenStudy (anonymous):

Find: lim x->0 sin2x/3x?

OpenStudy (anonymous):

I know the answer is 2/3, but I don't understand why.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

\[\lim_{x\to 0}\frac{\sin(x)}{x}=1\]right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[\frac{\sin(2x)}{3x}=\frac{\sin(2x)}{2x}\times \frac{2}{3}\]

OpenStudy (anonymous):

Oh.. that makes more sense than the explanation my teacher gave. Thanks a ton!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i wonder what your teacher said